two positive numbers are in the ratio 3:4, the sum is 3 time the first and 2 times the second is 68. find the smaller number
Answer by checkley79(3341)

LET THE FIRST NUMBER BE 3N
LET THE SECOND NUMBER BE 4N
3*3M+2*4M=68
9M+MN=68
17M=68
M=68/17
M=4 THIS IS THE MULTIPLIER FOR THE RATIOS.
3*4=12 ANS. FOR THE SMALLER NUMBER.
4*4=16 ANS. FOR THE LARGER NUMBER.
PROOF:
3/4=12/16
4*12=3*16
48=48
3*12+2*16=68
36+32=68
68=68